Circular No. 6163

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

DETECTION OF FAINT KUIPER BELT OBJECTS The Hubble Space Telescope Kuiper Belt Search Team (A. Cochran,

University of Texas; H. Levison and A. Stern, Southwest Research Institute; and M. Duncan, Queens University) reports: “We have found several candidate Kuiper-belt objects using deep HST/WFPC2 images obtained in the ecliptic near morning quadrature. These objects are quite faint (V ∼ 28), suggesting that they are comparable in size to comet 1P/Halley. The observations consist of 34 exposures of ∼ 10 min each of a single 4-square-arcmin field centered at α = 3^{h}41^{m}20^{s}, δ = +19^{o}34′.7 (equinox 2000.0); the exposures were obtained on a number of consecutive orbits during 1994 Aug. 21.7–21.9 and 22.7–23.0 UT. Each WF chip was handled separately so that no assumptions were made about relative orientation of the chips. The first step in our data reduction was to sum all of the images with no shifts to determine the positions of stars and galaxies; this median sum was subtracted from each image prior to further processing. The stellar and galactic images had a typical WFPC2 PSF, indicating that the telescope was pointing at the same field on each of the 34 images. After removal of stars and galaxies, we stacked the resultant images with a variety of shifts to simulate the drift rates of objects in the outer solar system. A typical Kuiper-belt orbit shifts by ∼ 150 pixels in the course of the 34 observations. So far, we have searched for objects in 44 viable Kuiper-belt orbits. These orbits all are in the 2:3 mean-motion resonance with Neptune, have eccentricities between 0.1 and 0.3 and inclinations < 3^{o}.3 to the ecliptic. In addition, we also searched 44 unrealistic orbits as a control. With our data-reduction procedure we have found a total of 244 candidate objects in our real orbits but only 185 in our control orbits. A χ^{2} test indicates that there is less than a 1-percent chance that these two samples are drawn from the same parent population. In addition, a comparison of the two flux distributions shows that candidates in the real orbits tend to be brighter than those in the control sample. A visual inspection of several of the brighter candidates confirms that they are most likely real. If our 59 excess candidates are indeed real members of the Kuiper belt, there must be ∼ 60 000 such objects per square degree, or at least a total of 10^{8} comets brighter than our limiting magnitude in the restricted range of orbits similar to the ones studied here.”

1995 April 17 Daniel W. E. Green